Solve the inequality
X²-|x|-12
________ is less than or equal to 0
|X|
1
Expert's answer
2020-08-12T18:24:25-0400

Linear Algebra

We need to solve the given inequality "\frac {x^2 - |x| - 12} {|x|}\le 0"


Solution:

Given inequality


"\frac {x^2 - |x| - 12} {|x|}\le 0"

Multiply with "|x|" on both sides of inequality, then we get

"x^2 - |x| - 12 \le 0 \space and \space |x|>0"


We can also write this as,

"|x|^2 - |x| - 12\le0 \space and \space |x|>0"

We can factorize this as,

"|x|^2 - 4|x| + 3|x| - 12 \le 0"

"|x| ( |x| - 4) + 3 (|x| - 4) \le 0"

"(|x|+ 3) (|x| - 4) \le 0"

"(|x|- (- 3)) (|x| - 4) \le 0"

We can write the |x| as,

"- 3\le |x| \le 4 \space and \space |x|>0"

But |x| should not be negative and greater than 0


So, the solution would be,

"0< |x| \le 4"